OpenStudy (anonymous):

I have been using every math formula I know to figure this out what am I doing wrong? I would really appreciate some help. thanks :) Kate currently has an account balance of $7,194.66. She opened the account 21 years ago with a deposit of $2,978.41. If the interest compounds daily, what is the interest rate on the account?

OpenStudy (anonymous):

Use \[\large A=P(1+R)^t\]

OpenStudy (anonymous):

In this case, since the compounding is daily, you have to convert 21 years to days for 't'

OpenStudy (anonymous):

alright thanks im gonna try to see if this works thanks!

OpenStudy (anonymous):

so would I divide 21 by 7?

OpenStudy (anonymous):

Number of days in 21 years = (21years) × (365.25days/year).

OpenStudy (anonymous):

Also remember that since this is being solved in terms of days, you need to multiply the R that you get by 365.25 to get it in terms of years (APR).

OpenStudy (anonymous):

okey dokey thanks !!! :)

OpenStudy (amistre64):

im wondering if "compound daily" is the same as "continuous"

OpenStudy (anonymous):

hmmm.... then I should use the formula A=Pe^rt

OpenStudy (amistre64):

Pert would be my idea \[A=Pe^{rt}\] \[ln(\frac{A}{P})=rt\] \[\frac{1}{t}ln(\frac{A}{P})=r\]

OpenStudy (anonymous):

ok im trying that thanks.

OpenStudy (zarkon):

"compound daily" is not the same as "continuous"...though it would prob be a good approximation

OpenStudy (anonymous):

ok so it isnt the same? im now confused.

OpenStudy (turingtest):

daily=once a day interest is compounded continuous=every instant interest is compounded

OpenStudy (anonymous):

ahhh ok i get it now thank you :)

OpenStudy (amistre64):

Daily gives me: .0X2 Continuous gives me: .0X19976.... so yeah, they are pretty close

OpenStudy (anonymous):

Continuous compounding is a good approximation, but it would be good algebra practice to actually do it with daily.

OpenStudy (anonymous):

I'd start with \[\large A=P(1+R)^T\] \[\large \rightarrow log(A/P)=T \cdot log(1+R)\] \[\large let \space X=log(A/P)/T \rightarrow R=10^X-1\] That R is the daily rate using T=7670.25 days, so multiply by 365.25 to get the APR.

OpenStudy (anonymous):

alright then thanks everyone!!! :)